OpenStudy (31356):

y-x=4 in standard form MEDAL REWARDED!!

OpenStudy (smileyxl3):

@DaBaller already gave you the formula

OpenStudy (31356):

But I need more help Smiley

OpenStudy (anonymous):

Standard form = Ax+By=C So all you need to do is switch the x and y around, but remember since it says -x, the negative goes with it, making it... -x+y=4

OpenStudy (smileyxl3):

Ax + By = C -x+y=4

OpenStudy (31356):

Thx DaBaller :)